# Flow Matching for Tabular Data Synthesis

### v1

@@ 1 Introduction
Government bodies, particularly national statistical offices (NSOs), are confronted with a pressing challenge: how to disseminate useful tabular data while preserving individual privacy. Tabular data, which are structured data with rows and columns representing real-world entities and their attributes, are essential for decision-making in economics, healthcare, social sciences, and finance (Little2024synthetic). However, because these datasets often contain sensitive attributes, such as financial conditions or medical diagnoses, privacy protection laws such as the EU’s General Data Protection Regulation (GDPR) and Indonesia’s Personal Data Protection (PDP) Act limit access to them (ruggles2021role). This tension between data utility and confidentiality has spurred interest in methods that can generate synthetic tabular data that preserve statistical properties while reducing the risk of disclosure (nowok2016synthpop1).
@@
Recent advances in deep generative models have provided a promising approach to tabular data synthesis. Deep learning techniques, such as Generative Adversarial Networks (goodfellow2014generative), Variational Autoencoders (kingma2013auto1), and more recently, diffusion models (dickstein2015deep; ho2020denoising), have demonstrated impressive capabilities in generating synthetic data of various forms, including images and text, as well as tabular data (xu2019modeling; kotelnikov2023tabddpm).
@@
Diffusion models, in particular, have emerged as a powerful option for generative tasks, including the synthesis of tabular data. TabDDPM (kotelnikov2023tabddpm), a diffusion model tailored for tabular data, and other models such as TabSyn (zhang2024mixedtype) have gained attention for their ability to model intricate dependencies between variables in tabular datasets by iteratively transforming noise into structured data. The main strength of these models lies in their ability to handle the high-dimensional structure of tabular data, where the relationships between variables are often nonlinear and complex.
@@
Another recent development is flow matching (FM) (lipman2023flowmatchinggenerativemodeling; albergo2023stochasticinterpolantsunifyingframework; liu2022rectifiedflowmarginalpreserving), which approximates the data distribution through a flow, i.e. through learning a velocity field that induces an ordinary differential equation transporting noise to data. Variational Flow Matching (VFM) (eijkelboom2024variationalflowmatchinggraph) offers a reinterpretation of flow matching as a form of variational inference over trajectories, providing an elegant solution to parameterise the flow towards any kind of distribution (e.g. also discrete or constrained). This method holds significant promise for tabular data synthesis, as it ensures that the generated output not only mirrors the statistical attributes of the original data but also captures patterns that traditional generative techniques might overlook. Beyond tabular data generation (guzman-cordero2025exponential), VFM has also seen recent success in fields like molecular generation (zaghen2025riemannianvariationalflowmatching; eijkelboom2025controlled; sakalyan2025modeling), image generation (maticsan2025purrception), and climate modeling (finn2025generative). Further discussion on related work can be found in Appendix A.
@@
Although recent advances in tabular diffusion models (kotelnikov2023tabddpm; zhang2024mixedtype) and FM (guzman-cordero2025exponential) have shown promise, existing studies remain limited in four ways. First, they have focused primarily on open source benchmark datasets, neglecting the unique challenges of census data as a real-world representation. Second, utility and privacy risk evaluation has been restricted to computer science metrics such as ML accuracy and distance to closest record (DCR), which (i) may not align with statistical needs,(ii) are computationally intensive, and (iii) can be hard to interpret for practitioners. Third, recent flow matching studies (eijkelboom2024variationalflowmatchinggraph; guzman-cordero2025exponential; eijkelboom2025controlled) have predominantly focused on optimal transport trajectories. The potential of variance-preserving trajectories - although theoretically established in the original FM formulation (lipman2023flowmatchinggenerativemodeling) and proven effective in diffusion models (ho2020denoising) — remains underexplored. Finally, the interaction between latent vs. direct representations and deterministic vs. stochastic dynamics in FM-based tabular synthesis remains unclear, leaving open questions about which configuration produces the best synthetic data.
@@
Contributions. We explore flow matching techniques for tabular data synthesis through four implementations: (1) learning the distribution of latent variables using regular flow matching, inspired by TabSyn (zhang2024mixedtype); (2) learning the tabular data distribution directly using variational flow matching; (3) a systematic comparison of interpolation schemes; and (4) implementing stochastic dynamics in VFM.
@@ 2.1 Flow Matching
Flow matching (FM) (lipman2023flowmatchinggenerativemodeling; liu2022flowstraightfastlearning; albergo2023stochasticinterpolantsunifyingframework) is a simulation-free generative modelling framework that learns a velocity field $v_{t}^{\theta}$ parameterised by $\theta\in\mathbb{R}^{p}$. This field defines a continuous transformation of a base distribution $p_{0}$ (typically a standard Gaussian) into a target distribution $p_{1}$ (empirical data) over a pseudo-time interval $t\in[0,1]$ through its induced ordinary differential equation.
@@ 2.2 Variational Flow Matching
Although flow matching is already an efficient and simulation free generative modelling framework, it can struggle with multimodal data or data of heterogeneous types (zhai2025vfpvariationalflowmatchingpolicy). To address this limitation, eijkelboom2024variationalflowmatchinggraph put forward variational flow matching (VFM), which reinterprets flow matching using the perspective of variational inference.
Unlike standard FM - which is primarily designed for continuous data - VFM allows a more flexible learning mechanism that accommodates a wider range of data types, including tabular data (guzman-cordero2025exponential). VFM learns an approximate vector field via the posterior distribution of the data using a variational distribution $q^{\theta}_{t}$:
@@
$\mathcal{L}_{\text{MF-VFM}}(\theta)$ $=\mathbb{E}_{t}\left[\text{KL}\left(p_{t}(x_{t})p_{t}(x_{1}\mid x_{t})\,\Big\|\,p_{t}(x_{t})q_{t}^{\theta}(x_{1}\mid x_{t})\right)\right]$ (2.6)
$=-\mathbb{E}_{t,x_{1}x_{t}}\left[\log q_{t}^{\theta}(x_{1}\mid x_{t})\right]+\text{const}$ (2.7)
$=-\mathbb{E}_{t,x_{1},x_{t}}\left[\sum_{d=1}^{D}\log q_{t}^{\theta}(x_{1}^{d}\mid x_{t})\right]+\text{const},$ (2.8)
where the final line assumes a mean-field approximation (which does not restrict expressivity as posited in the theoretical analysis of eijkelboom2024variationalflowmatchinggraph) factorising the variational posterior across dimensions. Having learned the posterior distribution, the vector field can be computed using equation 2.9:
@@
Beyond the deterministic ODE formulation, VFM also admits a stochastic counterpart that preserves the same marginal distribution $p_{t}(x)$ eijkelboom2024variationalflowmatchinggraph:
$\mathrm{d}x_{t}\;=\;\Big[u_{t}(x_{t})+\tfrac{g_{t}^{2}}{2}\,\nabla_{x_{t}}\log p_{t}(x_{t})\Big]\mathrm{d}t\;+\;g_{t}\,\mathrm{d}w_{t},$ (2.10)
where $g_{t}\!\geq\!0$ is a diffusion scheduler and $w_{t}$ is a standard Wiener process. The scheduler controls the stochasticity of the trajectory, but not the marginal path; Setting $g_{t}\!\equiv\!0$ recovers the ODE. We approximate this SDE using the variational distribution $q_{t}^{\theta}(x_{1}\mid x)$ as follows:
@@
where the drift and score are estimated as:
$v_{t}^{\theta}(x_{t})=\mathbb{E}_{q_{t}^{\theta}(x_{1}\mid x_{t})}[u_{t}(x_{t}\mid x_{1})],\qquad s_{t}^{\theta}(x_{t})=\mathbb{E}_{q_{t}^{\theta}(x_{1}\mid x_{t})}[\nabla_{x_{t}}\log p_{t}(x_{t}\mid x_{1})].$ (2.12)
@@
2.3 Trajectory Setting in Flow Matching
There are two commonly used trajectories to define the probability path in flow matching (lipman2023flowmatchinggenerativemodeling).
The first is the optimal transport (OT) trajectory (liu2022flowstraightfastlearning; tong2024improvinggeneralizingflowbasedgenerative), where the mean and standard deviation of the conditional distribution evolve linearly over time. This produces a deterministic vector field with a fixed direction, resulting in straight-line (conditional) paths connecting the noise and data distributions (liu2022flowstraightfastlearning). OT-based trajectories are computationally straightforward, efficient, and easy to parameterise due to their time-invariant structure, as shown in equation 2.13
$\begin{split}x_{t}&=tx_{1}+(1-(1-\sigma_{\text{min}})t)x_{0}\\ u_{t}(x_{t}\mid x_{1})&=\frac{x_{1}-(1-\sigma_{\text{min}})x_{t}}{1-(1-\sigma_{\text{min}})t},\end{split}$ (2.13)
The second type of trajectory is the variance-preserving (VP) path, derived from stochastic differential equations commonly used in diffusion models (lipman2023flowmatchinggenerativemodeling). In this approach, noise is gradually added to the data during the forward process, while the reverse process learns to de-noise using time-dependent vector fields. Unlike OT paths, VP diffusion trajectories evolve smoothly in both magnitude and direction over time. The parameters $\alpha_{t}$ and $\sigma_{t}$ are determined by cumulative functions of the noise scale function $\beta(t)=\beta_{\text{min}}+t(\beta_{\text{max}}-\beta_{\text{min}})$, as given in equation 2.14.
$\begin{split}\alpha_{t}&=e^{-\frac{1}{2}T(t)},\quad T(t)=\int_{0}^{1-t}\beta(s)ds,\quad\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}\\ u_{t}(x_{t}|x_{1})&=-\frac{\dot{T}(t)}{2}\left[\frac{e^{-T(t)}x_{t}-e^{-\frac{1}{2}T(t)}x_{1}}{1-e^{-T(t)}}\right]\end{split}$ (2.14)
The choice of trajectory – or interpolation scheme – significantly affects data synthesis performance, which we explore in this paper. Depending on the data structure or modelling objectives, either trajectory can be more suitable. Additional trajectory formulations are provided in Appendix E.1.
2.4 TabSyn: Latent Diffusion for Tabular Data
TabSyn (zhang2024mixedtype) represents a state-of-the-art generative approach for mixed-type tabular data. It integrates a transformer-based Variational Autoencoder (VAE) with a score-based diffusion model operating in latent space. The design of TabSyn addresses several key challenges in tabular data synthesis including heterogeneous feature types, complex inter-feature dependencies, and the inefficiency of diffusion processes when employed on high-dimensional structured data.
Within the latent space, noise is introduced through a forward process defined as: $z_{t}=z_{0}+\sigma(t)\varepsilon$, with $\sigma(t)=t$ under a linear noise schedule. The corresponding reverse process is an SDE:
$\mathrm{d}z_{t}=-2\sigma_{t}\dot{\sigma}_{t}\nabla_{z}\log p_{t}(z_{t})\,\mathrm{d}t+\sqrt{2\sigma_{t}\dot{\sigma}(t)}\,\mathrm{d}\omega_{t}.$ (2.15)
where $\omega_{t}$ denotes a Wiener process. The model is trained using denoising score matching, with a neural network $\varepsilon_{\theta}(z_{t},t)$ that estimates the scoring function $\nabla_{z_{t}}\log p_{t}(z_{t})=-\varepsilon_{\theta}(z_{t},t)/\sigma(t)$.
A study by  zhang2024mixedtype showed that TabSyn surpasses traditional GAN and VAE based models as well as the latest diffusion models like TabDDPM and STaSy, across a variety of datasets and evaluation metrics - including unvariate distribution and pairwise correlation preservation, and downstream ML performance. Furthermore, its latent-space formulation allows for missing value imputation via latent inpainting, achieving competitive results in terms of data quality, diversity, and privacy preservation.
@@ 3.1 Motivation and Design Axes
Flow Matching provides a general recipe for learning deterministic or stochastic transformations between probability distributions. Yet, its effectiveness in practice depends not on a single equation, but on a set of design decisions that determine how distributions are represented, interpolated, and evolved over time. When applied to tabular data synthesis, these decisions interact in particularly complex ways: mixed continuous-categorical features challenge smooth trajectories, privacy-sensitive settings impose computational limits, and statistical agencies often require explainable and reproducible models. This section therefore reformulates Flow Matching as a design space rather than a single algorithmic recipe, and places tabular synthesis at its centre.
@@
    Trajectory defines the interpolant between source and data distributions – commonly Optimal Transport (OT) or Variance Preserving (VP) paths – which control how signal and noise evolve over pseudo-time.
@@
    Dynamics determine whether the learned flow is integrated deterministically as an ODE or stochastically as an SDE, offering a continuum between precise transport and exploratory regularisation.
By systematically combining these axes, we focus on two complementary formulations. First, we introduce TabSynFlow, which learns deterministic ODE flows in latent space, substituting the diffusion process in TabSyn with a learned velocity field. We also evaluate TabbyFlow developed by guzman-cordero2025exponential which extends Variational Flow Matching directly to data space, supporting both numerical and categorical variables through hybrid Gaussian-categorical posteriors.
@@ 3.2 TabbyFlow: Variational Flow Matching in Data Spaces
TabbyFlow (guzman-cordero2025exponential) extends the Variational Flow Matching (VFM) framework of eijkelboom2024variationalflowmatchinggraph to structured tabular data containing both numerical and categorical features. This formulation enables direct modelling in the data space, avoiding the need for an intermediate latent representation, while explicitly accounting for heterogeneous variable types within the flow dynamics.
The TabbyFlow objective treats each feature type according to its statistical nature: continuous variables are modelled using a Gaussian distribution, while categorical variables are modelled using categorical distributions. Let $x=(x_{\text{num}},x_{\text{cat}})$ denote a sample comprising $D_{\text{num}}$ numerical and $D_{\text{cat}}$ categorical features. Each categorical feature is one-hot encoded, yielding total dimensionality
$D=D_{\text{num}}+\sum_{d=1}^{D_{\text{cat}}}K_{d},$ (3.1)
where $K_{d}$ is the number of categories for the $d$-th categorical variable.
@@
Following Theorem 3 of eijkelboom2024variationalflowmatchinggraph, we assume a mean-field variational posterior $q_{t}^{\theta}(x_{\text{num},1}\mid x_{t})$ with Gaussian structure parameterised by mean $\theta_{t}(x_{t})$ and variance $0.5A_{t}(x_{t})^{-2}$. The VFM loss for numerical variables is
@@
However, we found that using $0.5A_{t}(x_{t})^{-2}$ as variance yielded lower performance in our experiment. To mitigate this, we implemented a relaxed variance of $0.5A_{t}(x_{t})^{-1}$. This adjustment slightly inflates the variance, therefore acting as a mild regulariser. The empirical results can be seen in appendix D.1.
@@
During sampling, a noise sample $\tilde{x}_{0}$ is drawn from $p_{0}$, and the trained network predicts $\theta_{t}(x_{t})$ at each time step. Because TabbyFlow does not directly output a velocity field, it is recovered via
$v_{t}^{\theta}(x_{t})=\left(\dot{\alpha}_{t}-\frac{\dot{\sigma}_{t}}{\sigma_{t}}\alpha_{t}\right)\theta_{t}(x_{t})+\frac{\dot{\sigma}_{t}}{\sigma_{t}}x_{t},$ (3.5)
and, for the stochastic variant, the score function
$s_{t}^{\theta}(x_{t})=-\frac{x_{t}-\alpha_{t}\theta_{t}(x_{t})}{\sigma_{t}^{2}}.$ (3.6)
These components define the ODE or SDE dynamics used for synthesis (Algorithms 34 in Appendix B.2). TabbyFlow thus enables a unified treatment of numerical and categorical features within a single variational flow, generalising continuous-time generative modelling to structured tabular domains.

### v2

@@ 1 Introduction
Government bodies, particularly national statistical offices (NSOs), are confronted with a pressing challenge: how to disseminate useful tabular data while preserving individual privacy. Tabular data, which are structured data with rows and columns representing real-world entities and their attributes, are essential for decision-making in economics, healthcare, social sciences, and finance (Little et al. 2024). However, because these datasets often contain sensitive attributes, such as financial conditions or medical diagnoses, privacy protection laws such as the EU’s General Data Protection Regulation (GDPR) and Indonesia’s Personal Data Protection (PDP) Act limit access to them (Ruggles and Van Riper 2021). This tension between data utility and confidentiality has spurred interest in methods that can generate synthetic tabular data that preserve statistical properties while reducing the risk of disclosure (Nowok et al. 2016).
@@
Recent advances in deep generative models have provided a promising approach to tabular data synthesis. Deep learning techniques, such as Generative Adversarial Networks (Goodfellow et al. 2014), Variational Autoencoders (Kingma and Welling 2013), and more recently, diffusion models (Sohl-Dickstein et al. 2015; Ho et al. 2020), have demonstrated impressive capabilities in generating synthetic data of various forms, including images and text, as well as tabular data (Xu et al. 2019; Kotelnikov et al. 2023).
@@
Diffusion models, in particular, have emerged as a powerful option for generative tasks, including the synthesis of tabular data. TabDDPM (Kotelnikov et al. 2023), a diffusion model tailored for tabular data, and other models such as TabSyn (Zhang et al. 2024) have gained attention for their ability to model intricate dependencies between variables in tabular datasets by iteratively transforming noise into structured data. The main strength of these models lies in their ability to handle the high-dimensional structure of tabular data, where the relationships between variables are often nonlinear and complex.
@@
Another recent development is flow matching (FM) (Lipman et al. 2023; Albergo et al. 2023; Liu 2022), which approximates the data distribution through a flow, i.e. through learning a velocity field that induces an ordinary differential equation transporting noise to data. Variational Flow Matching (VFM) (Eijkelboom et al. 2024) offers a reinterpretation of flow matching as a form of variational inference over trajectories, providing an elegant solution to parameterise the flow towards any kind of distribution (e.g. also discrete or constrained). This method holds significant promise for tabular data synthesis, as it ensures that the generated output not only mirrors the statistical attributes of the original data but also captures patterns that traditional generative techniques might overlook. Beyond tabular data generation (Guzmán-Cordero et al. 2025), VFM has also seen recent success in fields like molecular generation (Zaghen et al. 2025b; Eijkelboom et al. 2025; Sakalyan et al. 2025), image generation (Matişan et al. 2025), and climate modeling (Finn et al. 2025). Further discussion on related work can be found in Appendix A.
@@
Although recent advances in tabular diffusion models (Kotelnikov et al. 2023; Zhang et al. 2024) and FM (Guzmán-Cordero et al. 2025) have shown promise, existing studies remain limited in four ways. First, they have focused primarily on open source benchmark datasets, neglecting the unique challenges of census data as a real-world representation. Second, utility and privacy risk evaluation has been restricted to computer science metrics such as ML accuracy and distance to closest record (DCR), which (i) may not align with statistical needs,(ii) are computationally intensive, and (iii) can be hard to interpret for practitioners. Third, recent flow matching studies (Eijkelboom et al. 2024; Guzmán-Cordero et al. 2025; Eijkelboom et al. 2025) have predominantly focused on optimal transport trajectories. The potential of variance-preserving trajectories - although theoretically established in the original FM formulation (Lipman et al. 2023) and proven effective in diffusion models (Ho et al. 2020) — remains underexplored. Finally, the interaction between latent vs. direct representations and deterministic vs. stochastic dynamics in FM-based tabular synthesis remains unclear, leaving open questions about which configuration produces the best synthetic data.
@@
Contributions. We explore flow matching techniques for tabular data synthesis through four implementations: (1) learning the distribution of latent variables using regular flow matching, inspired by TabSyn (Zhang et al. 2024); (2) learning the tabular data distribution directly using variational flow matching; (3) a systematic comparison of interpolation schemes; and (4) implementing stochastic dynamics in VFM.
@@ 2.1 Flow Matching
Flow matching (FM) (Lipman et al. 2023; Liu et al. 2022; Albergo et al. 2023) is a simulation-free generative modelling framework that learns a velocity field $v_{t}^{\theta}$ parameterised by $\theta\in\mathbb{R}^{p}$. This field defines a continuous transformation of a base distribution $p_{0}$ (typically a standard Gaussian) into a target distribution $p_{1}$ (empirical data) over a pseudo-time interval $t\in[0,1]$ through its induced ordinary differential equation.
@@ 2.2 Variational Flow Matching
Although flow matching is already an efficient and simulation free generative modelling framework, it can struggle with multimodal data or data of heterogeneous types (Zhai and Hao 2025). To address this limitation, Eijkelboom et al. (2024) put forward variational flow matching (VFM), which reinterprets flow matching using the perspective of variational inference.
Unlike standard FM - which is primarily designed for continuous data - VFM allows a more flexible learning mechanism that accommodates a wider range of data types, including tabular data (Guzmán-Cordero et al. 2025). VFM learns an approximate vector field via the posterior distribution of the data using a variational distribution $q^{\theta}_{t}$:
@@
$\mathcal{L}_{\text{MF-VFM}}(\theta)$ $=\mathbb{E}_{t}\left[\text{KL}\left(p_{t}(x_{t})p_{t}(x_{1}\mid x_{t})\,\Big\|\,p_{t}(x_{t})q_{t}^{\theta}(x_{1}\mid x_{t})\right)\right]=\mathbb{E}_{t,x_{1},x_{t}}\left[\sum_{d=1}^{D}\log q_{t}^{\theta}(x_{1}^{d}\mid x_{t})\right]+\text{const}$ (2.6)
VFM uses mean-field variational inference, i.e., $q_{t}^{\theta}(x_{1}\mid x_{t})=\prod_{d=1}^{D}q_{t}^{\theta}(x_{1}^{d}\mid x_{t})$, which corresponds to assuming conditional independence across dimensions under the variational posterior $q_{t}^{\theta}$ (not under the true posterior $p_{t}(x_{1}\mid x_{t})$). Following Eijkelboom et al. (2024), this choice yields a tractable objective and still enables recovering the true distribution at $t=1$. The mean field formulation also supports flexible variational distributions, particularly exponential families (Guzmán-Cordero et al. 2025) such as Gaussian or categorical distribution. Having learned the posterior distribution approximation, the vector field can be computed using equation 2.7:
@@
Beyond the deterministic ODE formulation, VFM also admits a stochastic counterpart that preserves the same marginal distribution $p_{t}(x)$ that can be approximated using variational distribution (Eijkelboom et al. 2024):
@@
where $g_{t}\!\geq\!0$ is a diffusion scheduler and $w_{t}$ is a standard Wiener process. The scheduler controls the stochasticity of the trajectory, but not the marginal path; Setting $g_{t}\!\equiv\!0$ recovers the ODE. The drift is calculated using equation 2.7 and score $s_{t}^{\theta}(x_{t})$ is estimated as:
$s_{t}^{\theta}(x_{t})=\mathbb{E}_{q_{t}^{\theta}(x_{1}\mid x_{t})}[\nabla_{x_{t}}\log p_{t}(x_{t}\mid x_{1})]=-\frac{x_{t}-\alpha_{t}\theta_{t}(x_{t})}{\sigma_{t}^{2}}.$ (2.9)
@@ 3.1 Motivation and Design Axes
Flow Matching provides a general recipe for learning deterministic or stochastic transformations between probability distributions. Yet, its effectiveness in practice depends not on a single equation, but on a set of design decisions that determine how distributions are represented, interpolated, and evolved over time. When applied to tabular data synthesis, these decisions interact in particularly complex ways. First, mixed continuous and categorical features require explicit representation choices, and moving to a continuous latent space can simplify the flow dynamics but does not remove the categorical challenge because discrete structure must still be captured and decoded. Second, privacy-sensitive settings impose computational limits, and statistical agencies often require explainable and reproducible model configurations. This section therefore reformulates Flow Matching as a design space rather than a single algorithmic recipe, and places tabular synthesis at its centre.
@@
    Trajectory defines the interpolant between source and data distributions – commonly Optimal Transport (OT) or Variance Preserving (VP) paths – which control how signal and noise evolve over time.
@@
    Dynamics determine whether the learned flow is integrated deterministically as an ODE or stochastically as an SDE, corresponding to a deterministic transport map versus a stochastic process.
By systematically combining these axes, we focus on two complementary formulations. First, we introduce TabSynFlow, which learns deterministic ODE flows in latent space, substituting the diffusion process in TabSyn with a learned velocity field. We also evaluate TabbyFlow developed by Guzmán-Cordero et al. (2025) which extends Variational Flow Matching directly to data space, supporting both numerical and categorical variables through hybrid Gaussian-categorical posteriors.
@@ 3.2 TabbyFlow: Variational Flow Matching in Data Spaces
TabbyFlow (Guzmán-Cordero et al. 2025) extends the Variational Flow Matching (VFM) framework of Eijkelboom et al. (2024) to structured tabular data containing both numerical and categorical features. This formulation enables direct modelling in the data space, avoiding the need for an intermediate latent representation, while explicitly accounting for heterogeneous variable types within the flow dynamics.
The TabbyFlow objective treats each feature type according to its statistical nature: continuous variables are modelled using a Gaussian distribution, while categorical variables are modelled using categorical distributions. Let $x=(x_{\text{num}},x_{\text{cat}})$ denote a sample comprising $D_{\text{num}}$ numerical and $D_{\text{cat}}$ categorical features. Each categorical feature is one-hot encoded, yielding total dimensionality $D=D_{\text{num}}+\sum_{d=1}^{D_{\text{cat}}}K_{d}$, where $K_{d}$ is the number of categories for the $d$-th categorical variable.
@@
Following Theorem 3 of Eijkelboom et al. (2024), we assume a mean-field variational posterior $q_{t}^{\theta}(x_{\text{num},1}\mid x_{t})$ with Gaussian structure parameterised by mean $\theta_{t}(x_{t})$ and variance $0.5A_{t}(x_{t})^{-2}$. The VFM loss for numerical variables is
@@
However, we found that using $0.5A_{t}(x_{t})^{-2}$ as variance yielded lower performance in our experiment. To mitigate this, we implemented a relaxed variance of $0.5A_{t}(x_{t})^{-1}$. This adjustment slightly inflates the variance, therefore acting as a mild regulariser. The empirical results can be seen in appendix G.1.
@@
During sampling, a noise sample $\tilde{x}_{0}$ is drawn from $p_{0}$, and the trained network predicts $\theta_{t}(x_{t})$ at each time step. Because TabbyFlow does not directly output a velocity field, it is recovered via equation 2.7, while the score function $s_{t}^{\theta}(x_{t})$ via equation 2.9. These components define the ODE or SDE dynamics used for synthesis (Algorithms 34 in Appendix C.2). TabbyFlow thus enables a unified treatment of numerical and categorical features within a single variational flow, generalising continuous-time generative modelling to structured tabular domains.